4) 15-P17, chap 15 problem 17 (a) A cyclical heat engine, operating between temperatures of 450° C and 150° C produces 4.00 MJ of work on a heat transfer of 5.00 MJ into the engine. How much heat transfer occurs to the environment? (b) What is unreasonable about the engine? (c) Which premise is unreasonable?
Answer #1

(a) heat transfer to environment = Q- W = (5 - 4) = 1 MJ

(b) maximum efficiency = 1 - Tc/Th

= 1 - (273 +150) / (273 + 450)

= 0.41

= 41%

efficiency of the engine = 4/5 * 100 = 80 %

since the efficiency of the engine is greater than the maximum efficiency it is not possible